• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 22:10:52 on August 4,2006

    Entry number 177460

    keyword=Setup Spectrometers for KIN#3

    Since MCC was sending us tune beam again to try to fix the beam position problems we've been having, we decided to pack it in for KIN# 10a and set up for KIN #3 in the event we get the beam back.

    The spectrometers are now set up and we will move the target when we get beam back. This has been a very frustrating shift.